# CaptainCook4D: A Dataset for Understanding Errors in Procedural Activities

### v3

@@ Abstract
Following step-by-step procedures is an essential component of various activities carried out by individuals in their daily lives. These procedures serve as a guiding framework that helps to achieve goals efficiently, whether it is assembling furniture or preparing a recipe. However, the complexity and duration of procedural activities inherently increase the likelihood of making errors. Understanding such procedural activities from a sequence of frames is a challenging task that demands an accurate interpretation of visual information and the ability to reason about the structure of the activity. To this end, we collect a new egocentric 4D dataset CaptainCook4D comprising 384 recordings (94.5 hours) of people performing recipes in real kitchen environments. This dataset consists of two distinct types of activities: one in which participants adhere to the provided recipe instructions and another in which they deviate and induce errors. We provide 5.3K step annotations and 10K fine-grained action annotations and benchmark the dataset for the following tasks: error recognition, multi-step localization and procedure learning111website: https://captaincook4d.github.io/captain-cook/.
@@ 2 Related Work
| ✓ | ✓ | EgoTV (Rishi Hazra, 2023) | Cooking / Simulated | ✓ | - | ✓ | Intentional | 168 | 2023 | |
@@ 5 Discussion, Limitations and Future Work
Discussion. We introduced a novel egocentric dataset for understanding errors in procedural activities. Our dataset consists of synchronized egocentric views, audio, and depth information specifically designed for tasks such as 3D activity analysis, Procedure Learning, Error Recognition, and more. While current methods have yielded promising outcomes, they continue to struggle to tackle these challenges adequately with satisfactory results, as demonstrated by our experimental assessment. This indicates the need for further exploration in this domain. Limitations. We aimed to capture deviations during procedural activities from an egocentric perspective. Since such data cannot be sourced from crowd-sourced platforms we captured participant data while performing procedural activities. By the nature of the problem, errors that occur when performing procedural activities are combinatorial and can have a compounding effect. Thus, our work has the following limitations: (1) For each activity, the errors captured and presented in the dataset form a subset of the whole combinatorial space; (2) Capturing 4D data in real kitchen environments posed logistical and equipment training challenges. As a result, we were compelled to limit the data collection to a specific geographic area. Future Work. Our work opens up several avenues for future work. First, an exciting direction is the extension of the dataset to include activities from other domains. The dataset can encompass a wider range of activities by incorporating tasks such as executing hardware-related activities (e.g., working with cars or computer parts). Second, the dataset can be used to compare and develop methods for solving tasks such as Few-Shot Error Recognition using visual/textual prompts, Semantic Role Labelling, Long Video Understanding, Procedure Planning, Reducing Errors, etc.
@@ A.1 Motivation
Error recognition in procedural activities has received significant traction, leading to the proposal of new datasets with errors [Wang et al., 2023, Ghoddoosian et al., 2023, Rishi Hazra, 2023, Schoonbeek et al., 2024]. Although they aim to identify errors in procedural activities, they focus on tasks related to assembly and disassembly. The activities involve objects with constant shapes and colors, which lack the desired characteristics. The absence of such specific video resources led us to curate a dataset (Fig. 8) embodying all our desired characteristics. By focusing on cooking activities with desired characteristics, our dataset can be used to develop easily transferable algorithms for other sensitive domains, such as medicine and chemistry.
@@ A.2 Extended Related Work
Recently, [Seminara et al., 2024] proposed a method for supervised procedure learning on the proposed dataset. A complete survey of all the relevant tasks is outside the scope of the paper; thus, we provide a brief review of procedure understanding tasks that are of particular interest to the proposed dataset such as Error Recognition[Soran et al., 2015, Boyang Wan et al., 2021], Multi-Step Localization and Self-Supervised Procedure Learning [Dwibedi et al., 2019, Elhamifar and Huynh, 2020, L. Logeswaran et al., 2023], Video Summarization [Narasimhan, Medhini et al., 2022], Temporal Action Segmentation [Fangqiu Yi et al., 2021, Li et al., 2019, Abu Farha et al., 2019, Chang et al., 2019, Fried et al., 2020], Object State Change Detection [Souček, Tomáš et al., 2022], Action Localization [Zhukov et al., 2020, Singh et al., 2017], Adverb Recognition [D. Moltisanti et al., 2023, 2023], Task Verification (Sequence Verification [Yichen Qian et al., 2021]) [Rishi Hazra, 2023], Long Video Understanding [Islam, Md Mohaiminul and Bertasius, Gedas, 2022, Honglu Zhou et al., 2023, Yue Zhao et al., 2022], Key-Step Localization [Nikita Dvornik et al., 2023, Malmaud et al., 2015, Hendricks et al., 2018], Procedure Planning [Henghui Zhao et al., 2022] (Goal-Step Inference [Yue Yang et al., 2021, Kuehne et al., 2014b, Pei et al., 2011]), Self-supervised procedural knowledge extraction [Nevatia et al., 2003] (Visual Transformation Telling [Xin Hong et al., 2023]), Sequence-to-sequence alignment [Nikita Dvornik et al., 2021, Haresh et al., 2021, Behrmann, Nadine et al., 2022, Bojanowski et al., 2015, Tengda Han et al., 2022, Junnan Li et al., 2021, Junru Wu et al., 2022, Sarkar, Pritam and Etemad, Ali, 2022, Jiahao Zhang et al., 2023, Sun, Yuchong et al., 2022, Yizhen Chen et al., 2023, Sixun Dong et al., 2023, Weizhe Liu et al., 2021] (Audio, Video synchronization [Morgado et al., 2020, Bowen Shi et al., 2022, Po-Yao Huang et al., 2022, Sihan Chen et al., 2023]), Scene Graph Anticipation [Peddi et al., 2024], Temporal Adaptation, Semantic Role Labelling [Arka Sadhu et al., 2021, Bojanowski et al., 2014, Huang et al., 2016], Procedure Learning [Zhou et al., 2018, Kaelbling, 1993, Madeline Chantry Schiappa and Y. Rawat, 2022, Zhao, Xueliang et al., 2022], Action Anticipation in Procedural Videos [Sener and Yao, 2019, Piergiovanni et al., 2020, Alexandros Stergiou and D. Damen, 2022].

### v4

@@ Abstract
Following step-by-step procedures is an essential component of various activities carried out by individuals in their daily lives. These procedures serve as a guiding framework that helps to achieve goals efficiently, whether it is assembling furniture or preparing a recipe. However, the complexity and duration of procedural activities inherently increase the likelihood of making errors. Understanding such procedural activities from a sequence of frames is a challenging task that demands an accurate interpretation of visual information and the ability to reason about the structure of the activity. To this end, we collect a new egocentric 4D dataset CaptainCook4D comprising 384 recordings (94.5 hours) of people performing recipes in real kitchen environments. This dataset consists of two distinct types of activities: one in which participants adhere to the provided recipe instructions and another in which they deviate and induce errors. We provide 5.3K step annotations and 10K fine-grained action annotations and benchmark the dataset for the following tasks: error recognition, multi-step localization and procedure learning111website: https://rohithpeddi.github.io/#/captaincook.
@@ 2 Related Work
| ✓ | ✓ | EgoTV (Hazra et al., 2023) | Cooking / Simulated | ✓ | - | ✓ | Intentional | 168 | 2023 | |
@@ 5 Discussion, Limitations and Future Work
Discussion. We introduced a novel egocentric dataset for understanding errors in procedural activities. Our dataset consists of synchronized egocentric views, audio, and depth information specifically designed for tasks such as 3D activity analysis, Procedure Learning, Error Recognition, and more. While current methods have yielded promising outcomes, they continue to struggle to tackle these challenges adequately with satisfactory results, as demonstrated by our experimental assessment. This indicates the need for further exploration in this domain. Limitations. We aimed to capture deviations during procedural activities from an egocentric perspective. Since such data cannot be sourced from crowd-sourced platforms we captured participant data while performing procedural activities. By the nature of the problem, errors that occur when performing procedural activities are combinatorial and can have a compounding effect. Thus, our work has the following limitations: (1) For each activity, the errors captured and presented in the dataset form a subset of the whole combinatorial space; (2) Capturing 4D data in real kitchen environments posed logistical and equipment training challenges. As a result, we were compelled to limit the data collection to a specific geographic area. Future Work. Our work opens up several avenues for future work. First, an exciting direction is the extension of the dataset to include activities from other domains. The dataset can encompass a wider range of activities by incorporating tasks such as executing hardware-related activities (e.g., working with cars or computer parts). Second, the dataset can be used to compare and develop methods for solving tasks such as Object State Change Detection (Souček, Tomáš et al., 2022), Action Localization (Zhukov et al., 2020; Singh et al., 2017), Adverb Recognition (D. Moltisanti et al., 2023, 2023), Task Verification (Sequence Verification (Yichen Qian et al., 2021)) (Hazra et al., 2023), Long Video Understanding (Islam, Md Mohaiminul and Bertasius, Gedas, 2022; Honglu Zhou et al., 2023; Yue Zhao et al., 2022), Robust Learning (Peddi et al., 2022; Peddi and Gogate, 2022), Key-Step Localization (Nikita Dvornik et al., 2023; Malmaud et al., 2015; Hendricks et al., 2018), Procedure Planning (Henghui Zhao et al., 2022), Self-supervised procedural knowledge extraction (Nevatia et al., 2003) (Visual Transformation Telling (Xin Hong et al., 2023)), Scene Graph Generation (Peddi et al., 2024; Rodin et al., 2023), Scene Graph Anticipation (Peddi et al., 2025), Temporal Adaptation, Semantic Role Labelling (Arka Sadhu et al., 2021; Bojanowski et al., 2014; Huang et al., 2016), Action Anticipation in Procedural Videos (Sener and Yao, 2019; Piergiovanni et al., 2020; Alexandros Stergiou and D. Damen, 2022)
@@ A.1 Motivation
Error recognition in procedural activities has received significant traction, leading to the proposal of new datasets with errors [Wang et al., 2023, Ghoddoosian et al., 2023, Hazra et al., 2023, Schoonbeek et al., 2024]. Although they aim to identify errors in procedural activities, they focus on tasks related to assembly and disassembly. The activities involve objects with constant shapes and colors, which lack the desired characteristics. The absence of such specific video resources led us to curate a dataset (Fig. 8) embodying all our desired characteristics. By focusing on cooking activities with desired characteristics, our dataset can be used to develop easily transferable algorithms for other sensitive domains, such as medicine and chemistry.
@@ A.2 Extended Related Work
Recently, [Seminara et al., 2024] proposed a method for supervised procedure learning on the proposed dataset. A complete survey of all the relevant tasks is outside the scope of the paper; thus, we provide a brief review of procedure understanding tasks that are of particular interest to the proposed dataset such as Error Recognition[Soran et al., 2015, Boyang Wan et al., 2021], Multi-Step Localization and Self-Supervised Procedure Learning [Dwibedi et al., 2019, Elhamifar and Huynh, 2020, L. Logeswaran et al., 2023], Video Summarization [Narasimhan, Medhini et al., 2022], Temporal Action Segmentation [Fangqiu Yi et al., 2021, Li et al., 2019, Abu Farha et al., 2019, Chang et al., 2019, Fried et al., 2020], Object State Change Detection [Souček, Tomáš et al., 2022], Action Localization [Zhukov et al., 2020, Singh et al., 2017], Adverb Recognition [D. Moltisanti et al., 2023, 2023], Task Verification (Sequence Verification [Yichen Qian et al., 2021]) [Hazra et al., 2023], Long Video Understanding [Islam, Md Mohaiminul and Bertasius, Gedas, 2022, Honglu Zhou et al., 2023, Yue Zhao et al., 2022], Key-Step Localization [Nikita Dvornik et al., 2023, Malmaud et al., 2015, Hendricks et al., 2018], Procedure Planning [Henghui Zhao et al., 2022] (Goal-Step Inference [Yue Yang et al., 2021, Kuehne et al., 2014b, Pei et al., 2011]), Self-supervised procedural knowledge extraction [Nevatia et al., 2003] (Visual Transformation Telling [Xin Hong et al., 2023]), Sequence-to-sequence alignment [Nikita Dvornik et al., 2021, Haresh et al., 2021, Behrmann, Nadine et al., 2022, Bojanowski et al., 2015, Tengda Han et al., 2022, Junnan Li et al., 2021, Junru Wu et al., 2022, Sarkar, Pritam and Etemad, Ali, 2022, Jiahao Zhang et al., 2023, Sun, Yuchong et al., 2022, Yizhen Chen et al., 2023, Sixun Dong et al., 2023, Weizhe Liu et al., 2021] (Audio, Video synchronization [Morgado et al., 2020, Bowen Shi et al., 2022, Po-Yao Huang et al., 2022, Sihan Chen et al., 2023]), Scene Graph Anticipation [peddi2024scene], Temporal Adaptation, Semantic Role Labelling [Arka Sadhu et al., 2021, Bojanowski et al., 2014, Huang et al., 2016], Procedure Learning [Zhou et al., 2018, Kaelbling, 1993, Madeline Chantry Schiappa and Y. Rawat, 2022, Zhao, Xueliang et al., 2022], Action Anticipation in Procedural Videos [Sener and Yao, 2019, Piergiovanni et al., 2020, Alexandros Stergiou and D. Damen, 2022].